# Unlocking Multi-Dimensional Integration with Quantum Adaptive Importance Sampling

### v1

@@ Abstract
We introduce a quantum algorithm that performs Quantum Adaptive Importance Sampling (QAIS) for Monte Carlo integration of multidimensional functions, targeting in particular the computational challenges of high-energy physics. In this domain, the fundamental ingredients for theoretical predictions such as multiloop Feynman diagrams and the phase-space require evaluating high-dimensional integrals that are computationally demanding due to divergences and complex mathematical structures. The established method of Adaptive Importance Sampling, as implemented in tools like VEGAS, uses a grid-based approach that is iteratively refined in a separable way, per dimension. This separable approach efficiently suppresses the exponentially growing grid-handling computational cost, but also introduces performance drawbacks whenever strong inter-variable correlations are present. To utilize sampling resources more efficiently, QAIS exploits the exponentially large Hilbert space of a Parameterised Quantum Circuit (PQC) to manipulate a non-separable Probability Density Function (PDF) defined on a multidimensional grid. In this setting, entanglement within the PQC captures the correlations and intricacies of the target integrand’s structure. Then, the PQC strategically allocates samples across the multidimensional grid, focusing the samples in the small subspace where the important structures of the target integrand lie, and thus generates very precise integral estimations. As an application, we look at a very sharply peaked loop Feynman integral and at multi-modal benchmark integrals.
preprint: June 24, 2025
@@ 1 Introduction
These issues make brute-force Monte Carlo (MC) integration computationally prohibitive, and in some cases, entirely impractical, motivating the development of sophisticated variance reduction techniques such as Importance Sampling (IS). The most notable example is the VEGAS Lepage:1977sw ; Lepage:123074 ; Lepage:2020tgj algorithm, which performs Adaptive Importance Sampling with a hyper-rectangular grid, and has become the common standard. Specifically, VEGAS operates by iteratively adapting a piecewise-constant grid that concentrates samples in regions of high variance, dramatically reducing statistical uncertainties for a given budget of function evaluations. VEGAS and its variants, are assimilated into mainstream libraries (e.g. Cuba Hahn:2004fe , MadGraph Alwall:2011uj , and Sherpa Gleisberg:2008ta among others). More recently, integrators based on Machine Learning have emerged Bothmann:2020ywa ; Winterhalder:2021ngy , as well as VEGAS GPU-based optimizations to reduce runtime Carrazza:2020rdn .
@@
As classical integrators approach their practical limits, attention has shifted towards quantum computing, driven by theoretical advances in the field through quantum algorithms such as Shor’s integer factoring Shor_1997 , Grover’s unstructured search grover1996fastquantummechanicalalgorithm , and the Harrow-Hassidim-Lloyd quantum linear solver Harrow_2009 , as well as advancements in the quantum hardware. Quantum algorithms for MC integration have been developed recently deLejarza:2023qxk ; deLejarza:2024pgk ; deLejarza:2024scm ; Rodrigo:2024say ; Agliardi:2022ghn ; Williams:2025hza ; Cruz-Martinez:2023vgs . The majority of these approaches exploit the quadratic speedup offered by the fault-tolerant algorithm Quantum Amplitude Estimation (QAE) Brassard_2002 ; Montanaro_2015 .
@@
In this paper, we leverage one of the key strengths of quantum computers, their ability for sampling from probability distributions. Notably, one of the few demonstrations of quantum supremacy Arute:2019zxq , where a quantum device outperforms any classical counterpart, was fundamentally a sampling problem. This experiment highlighted the potential of quantum computers to surpass classical systems in specific tasks. With the current surge of generative modeling as a cornerstone of classical machine learning, and sampling being one of the strongest features of quantum devices, the intersection of these fields has attracted significant attention within the quantum computing community Romero2021 ; Paine2021 ; Zhu2022 ; Kyriienko2024 ; kasture2022protocols ; rudolph2023trainability . Notable examples include Quantum Generative Adversarial Networks (QGANs) Lloyd2018 ; Dallaire-Demers2018 ; Zoufal2019 ; Huang2021PRAppl , Quantum Boltzmann Machines (QBMs) Zoufal2021 ; Coopmans2024 , and Quantum Circuit Born Machines (QCBMs) Liu2018 ; Benedetti2019npj ; Benedetti_2020 ; Kiss2022 ; hibat2024framework . Among other applications, quantum generative modeling has shown particular potential across a wide range of problems within High-Energy Physics (HEP) Delgado:2022tpc ; Delgado:2023ofr ; Delgado:2024vne ; Bermot:2023kvh ; Tuysuz:2024hyl ; Rudolph:2023iqf ; deLejarza:2025upd , alongside other promising quantum machine learning approaches that have also been explored lejarza ; deLejarza:2022vhe ; Belis:2023atb ; Schuhmacher:2023pro ; Belis:2024guf ; Belis:2021zqi ; Casals:2025cuc .
@@ 2 Monte Carlo integration, Importance Sampling and VEGAS
so that with this transformation, a uniform draw in $y$-space concentrates samples in regions where the $\Delta{x}_{i}$ are the smallest. Thus, the objective becomes to use finer widths, where $|f({x})|$ is largest.
@@ 3.1 Constructing the Grid with a Parametrized Quantum Circuit
Equation (21) defines a piecewise‐uniform in the integration domain $\Omega$. Because a $Z$‐basis measurement samples exactly one of the $2^{n}$ basis states, the shot record $\{k_{(j_{1},\dots,j_{d})}\}$, follows a multinomial distribution with probabilities $p_{(j_{1},\dots,j_{d})}=|c_{(j_{1},\dots,j_{d})}|^{2}$. For a fixed PQC at $N$ measurements, the maximum‐likelihood estimator is simply the empirical frequency, ${\hat{p}_{(j_{1},\dots,j_{d})}}=k_{(j_{1},\dots,j_{d})}/N$. Thus, the BEE construction provides a direct way to construct the full $2^{n}$-cell grid, while the resolution along any dimension can be refined simply by adding qubits to that dimension.
@@ 3.2 Quantum Importance Sampling Statistical Framework
Based on these definitions, the total estimator for the integral over $\Omega$ is:
@@
Formally, the estimator of Eq. (27) can be re-expressed as:
$\hat{I}^{\rm(QAIS)}_{N}=\frac{1}{N}\sum_{i=1}^{2^{n}}w^{(i)}\sum_{j=1}^{N_{i}}f({\bf x}^{(i)}_{j})=\sum_{i=1}^{2^{n}}I_{i}~{}.$ (29)
After performing $N$ measurements, we observe a subset of the Hilbert space elements or grid cells, $\Omega^{-}\subset\Omega$. The remaining states (or grid cells), $\Omega\setminus\Omega^{-}$, are effectively unseen. The missing mass contribution from $\Omega\setminus\Omega^{-}$ introduces a bias to the integral estimation, leading to a consistent underestimation. This bias can be explicitly evaluated as
$\text{bias}=\mathbb{E}\left[\hat{I}_{\text{obs}}\right]-I=-\sum_{i\in\Omega\setminus\Omega^{-}}I_{i}<0~{},$ (30)
@@ 3.3 Debiasing Strategy using a Tiling Algorithm
$\tilde{I}=\sum_{i\in\Omega_{I}}w^{(i)}\sum_{j=1}^{N_{i}}f({\bf x}^{(i)}_{j})+\sum_{k\in\Omega_{N-I}}\tilde{w}^{(k)}\sum_{j=1}^{\tilde{N}_{k}}f({\bf x}^{(k)}_{j})~{},$ (31)
where the adjusted weights $\tilde{w}^{(k)}$ take into account the shapes of the new enlarged hyper-rectangular domains and $\tilde{N}_{k}$ refers to the number of random samples drawn within each enlarged domain $\tilde{\Omega}_{k}$. Through this construction, we explicitly include all previously neglected states, eliminating missing regions and ensuring that the estimator is unbiased.
In practice, for constructing the non-Important Region, we employ a Tiling algorithm. This algorithm generates a small number of contiguous, non-overlapping hyper-rectangles that collectively cover every previously unobserved grid cell, a process that corresponds to a full tiling of the $d$-dimensional integration domain. For the construction of such an algorithm, it is important to have a computational cost that scales polynomially with the number of observed states $M$, the dimension of the integration space $d$, and the number of qubits used for discretization $n$. It is an explicit and crucial constraint for the Tiling algorithm not to scale with the number of unobserved, Non-Important grid cells, since, in this framework, this is an exponential sized quantity.
@@
The details of the Tiling algorithm can be found in Appendix A. The procedure works by generating greedy expansions between blocks of different dimensions. By leveraging this approach, the algorithm skips large blocks of lower-ordered dimensions, when making expansions in a higher ordered dimension. In this way, it efficiently tiles arbitrarily sized gaps in the $d$-dimensional grid with hyper-rectangles, using at most $12(d-1)+$ 1 intervals per gap. This bound on the number of intervals is the key factor behind the suitable scaling of the Tiling algorithm. After accounting for all operations, even the arithmetic operations, the tile generation executes in time $\mathcal{O}(M\,d^{3}\,n^{2})$. This, compared to the state-sorting cost $\mathcal{O}(n\,M\log M)$, is typically negligible. Therefore, the asymptotic overhead coming from the tiling procedure is essentially $\mathcal{O}(M\log M)$.
@@ 3.4 Structure and Optimization of the Parametrized Quantum Circuit
To convert the continuous function $f$, into a discrete PDF or a Probability Mass Function (PMF) suitable for computing the discretized KL-Divergence, we proceed as follows. We first compute the function at $N_{i}$ representative points $\mathbf{x}_{i}$} within each observed grid cell $\Omega^{(i)}$, accounting in total to $N=\sum_{i=1}^{M}N_{i}$ function evaluations. Because the grid is uniform, every cell has the same volume, so common factors cancel in the normalization. Therefore, we construct the PMF by taking the simplified expressions, using $P(\Omega^{(i)})=f(\Omega^{(i)})/Z$ which is the probability of each grid cell, where $f(\Omega^{(i)})=1/N_{i}\sum_{i=1}^{N_{i}}f(\mathbf{x}_{i})$ is the within-cell function’s average and $Z=\sum_{i=1}^{M}f(\Omega^{(i)})$ the normalization constant. Ideally, for a better precision in the target PMF, the cell-wise sample count, $N_{i}$ would be updated after each run, by assigning more sample to high probability cells. This, combined with the monitoring of the within-cell variance in high impact region, can provide a criterion that indicates whether the grid resolution must be increased by adding more qubits. However, in this work, for robustness and simplicity, we use a fixed $N_{i}$ per cell and cache all evaluations, while pre-defining the grid’s size.
@@
Figure 2: Quantum Circuit architecture used for training. The figure features the general structure with the layer definition (top) as well as an explicit $5$ qubit example (bottom).
@@ 3.5 Absence of phantom peaks
where $\mathbf{r}_{0}=(0.23,0.23)$ and $\mathbf{r}_{1}=(0.74,0.74)$. Because the integral can be computed in closed form, it provides a clear benchmark for validating numerical methods. A normalization constant is chosen via analytical integration, so that the integral evaluates to $1$.
@@
To demonstrate that QAIS is unbiased, we conducted a series of nested MC experiments. Because the integral is analytically normalized to 1, this exact value serves as the ground truth for validating our numerical estimates. For each experiment, we performed $1000$ independent integration runs, each with a fixed proposal PDF from the same trained PQC. We then repeated these experiments with different shot counts: $1000$, $5000$ and $10000$ samples. The results are shown in Fig. 4. We see that the average integral value oscillates around the true value, improving accuracy progressively, and does not exhibit a systematic preference toward over or underestimation when the number of samples is varied. Additionally, the average QAIS standard deviation is presented. It shall be noted that estimates that deviate significantly from the mean integral value, always exhibit inflated standard deviation. Thus, QAIS produced stable accurate and unbiased estimates without the need to sample all $2^{n}$ possible states.
@@
where $\mathbf{r}=(0.5,0.5)$. Figure 5 illustrates the results for both methods. The exponential factor puts almost the entire weight of the integrand on a thin ring of radius $0.35$ centered at $\mathbf{r}$, while in the rest of the integration domain, it is exponentially suppressed and effectively negligible.
@@ 4.1 One-loop pentagon Feynman integral in the Loop-Tree Duality
The QAIS results are presented in Fig. 7 and Fig. 8. The training of the PQC was carried out with COBYLA, the Pennylane’s lightning.qubit noiseless state‑vector simulator, and different Anzätze with different numbers of qubits per dimension. It is worth noting that, in the smallest configuration we tested, i.e. the 16 qubits distributed as $(8,4,4)$ across the three dimensions and $2\times 10^{4}$ iterations, the Ansatz sequence $U^{(Z)}\to U_{3}$ with $144$ parameters reaches $D_{\rm KL}<0.5$ very fast and then saturates at $D_{\rm KL}\approx 0.27$. The results presented in Fig. 7 are obtained using the Ansatz $U^{(Z)}\to U_{3}\to U^{(X)}\to U_{3}$, which performed best given the constraints we imposed on the number of iterations. In this case, the number of parameters ranges from $336$ to $456$ for $16$ to $19$ qubits, respectively, and the number of iterations ranges from $2\times 10^{4}$ to $3\times 10^{4}$. The $D_{\rm KL}$ ranged from $0.09$ to $0.14$.
For the HEA, with the current number of iterations for the $(8,4,4)$ qubit case, we tested between $2$ to $8$ layers, corresponding to $96$ to $384$ parameters, with different random initializations. The results obtained were at least one order of magnitude larger in the KL divergence ($D_{\rm KL}>1$), indicating a significant discrepancy. It should be noted that the optimization that starts from a uniform superposition has initial $D_{\rm KL}\approx 6.2$.
@@
Figure 8: Histograms of the values of the 100 independent integration runs for the one-loop pentagon Feynman diagram with parameters in Eq. (40). The dashed vertical lines, correspond to $1\sigma$, $2\sigma$, and $3\sigma$ standard deviations. The reference value comes by the comparison with the classical numerical integration Buchta:2015wna using VEGAS, $I=-1.24027(16)\cdot 10^{-13}$.

### v2

@@ Abstract
We introduce a quantum algorithm that performs Quantum Adaptive Importance Sampling (QAIS) for Monte Carlo integration of multidimensional functions, targeting in particular the computational challenges of high-energy physics. In this domain, the fundamental ingredients for theoretical predictions such as multiloop Feynman diagrams and the phase-space require evaluating high-dimensional integrals that are computationally demanding due to divergences and complex mathematical structures. The established method of Adaptive Importance Sampling, as implemented in tools like VEGAS, uses a grid-based approach that is iteratively refined in a separable way, per dimension. This separable approach efficiently suppresses the exponentially growing grid-handling computational cost, but also introduces performance drawbacks whenever strong inter-variable correlations are present. To utilize sampling resources more efficiently, QAIS exploits the exponentially large Hilbert space of a Parameterised Quantum Circuit (PQC) to manipulate a non-separable Probability Density Function (PDF) defined on a multidimensional grid. In this setting, entanglement within the PQC captures the correlations and intricacies of the target integrand’s structure. Performing measurements on the PQC determines the sample allocation across the multidimensional grid. This focuses samples in the small subspace where the important structures of the target integrand lie, and thus generates very precise integral estimations. As an application, we look at a very sharply peaked loop Feynman integral and at multi-modal benchmark integrals.
preprint: July 25, 2025
@@ 1 Introduction
These issues make brute-force Monte Carlo (MC) integration computationally prohibitive, and in some cases, entirely impractical, motivating the development of sophisticated variance reduction techniques such as Importance Sampling (IS). The most notable example is the VEGAS Lepage:1977sw ; Lepage:123074 ; Lepage:2020tgj algorithm, which performs Adaptive Importance Sampling with a hyper-rectangular grid, and has become the common standard. Specifically, VEGAS operates by iteratively adapting a piecewise-constant grid that concentrates samples in regions of high variance, dramatically reducing statistical uncertainties for a given budget of function evaluations. VEGAS and its variants, are assimilated into mainstream libraries (e.g. Cuba Hahn:2004fe , MadGraph Alwall:2011uj , and Sherpa Gleisberg:2008ta among others). More recently, integrators based on Machine Learning have emerged Gao_2020 ; Gao_2020_2 ; Bothmann:2020ywa ; Winterhalder:2021ngy ; Heimel_2023 ; Heimel_2024 ; Heimel_2025 , as well as VEGAS GPU-based optimizations to reduce runtime Carrazza:2020rdn .
@@
As classical integrators approach their practical limits, attention has shifted towards quantum computing, driven by theoretical advances in the field through quantum algorithms such as Shor’s integer factoring Shor_1997 , Grover’s unstructured search grover1996fastquantummechanicalalgorithm , and the Harrow-Hassidim-Lloyd quantum linear solver Harrow_2009 , as well as advancements in the quantum hardware. Quantum algorithms for accelerating MC methods, mainly in numerical integration have been developed recently deLejarza:2023qxk ; deLejarza:2024pgk ; deLejarza:2024scm ; Rodrigo:2024say ; Agliardi:2022ghn ; Williams:2025hza ; Cruz-Martinez:2023vgs ; akhalwaya2023modularenginequantummonte ; Plekhanov_2022 ; Layden_2023 ; Mazzola_2021 . The majority of the integration approaches exploit the quadratic speedup offered by the fault-tolerant algorithm Quantum Amplitude Estimation (QAE) Brassard_2002 ; Montanaro_2015 .
@@
In this paper, we leverage one of the key strengths of quantum computers, their ability for sampling from probability distributions. Notably, one of the few demonstrations of quantum supremacy Arute:2019zxq , where a quantum device outperforms any classical counterpart, was fundamentally a sampling problem. This experiment highlighted the potential of quantum computers to surpass classical systems in specific tasks. With the current surge of generative modeling as a cornerstone of classical machine learning, and sampling being one of the strongest features of quantum devices, the intersection of these fields has attracted significant attention within the quantum computing community Romero2021 ; Paine2021 ; Zhu2022 ; Kyriienko2024 ; kasture2022protocols ; rudolph2023trainability . Notable examples include Quantum Generative Adversarial Networks (QGANs) Lloyd2018 ; Dallaire-Demers2018 ; Zoufal2019 ; Huang2021PRAppl , Quantum Boltzmann Machines (QBMs) Zoufal2021 ; Coopmans2024 , and Quantum Circuit Born Machines (QCBMs) Liu2018 ; Benedetti2019npj ; Benedetti_2020 ; Kiss2022 ; hibat2024framework . Among other applications, quantum generative modeling has shown particular potential across a wide range of problems within High-Energy Physics (HEP) Delgado:2022tpc ; Delgado:2023ofr ; Delgado:2024vne ; Bermot:2023kvh ; Tuysuz:2024hyl ; deLejarza:2025upd , alongside other promising quantum machine learning approaches that have also been explored lejarza ; deLejarza:2022vhe ; Belis:2023atb ; Schuhmacher:2023pro ; Belis:2024guf ; Belis:2021zqi ; Casals:2025cuc .
@@ 2 Monte Carlo integration, Importance Sampling and VEGAS
[…]
@@ 3.1 Constructing the Grid with a Parametrized Quantum Circuit
Equation (21) defines a piecewise‐uniform PDF in the integration domain $\Omega$. Because a $Z$‐basis measurement samples exactly one of the $2^{n}$ basis states, the shot record $\{k_{(j_{1},\dots,j_{d})}\}$, follows a multinomial distribution with probabilities $p_{(j_{1},\dots,j_{d})}=|c_{(j_{1},\dots,j_{d})}|^{2}$. For a fixed PQC at $N$ measurements, the maximum‐likelihood estimator is simply the empirical frequency, ${\hat{p}_{(j_{1},\dots,j_{d})}}=k_{(j_{1},\dots,j_{d})}/N$. Thus, the BEE construction provides a direct way to construct the full $2^{n}$-cell grid, while the resolution along any dimension can be refined simply by adding qubits to that dimension.
@@ 3.2 Quantum Importance Sampling Statistical Framework
Based on these notation definitions and by considering the form of the IS estimator of Eq. (5) and the form of the proposal PDF of Eq. (21), the total estimator for the integral over $\Omega$ is:
@@
After performing $N$ measurements, we observe a subset of the Hilbert space elements or grid cells, $\Omega^{-}\subset\Omega$. The remaining states (or grid cells), $\Omega\setminus\Omega^{-}$, are effectively unseen. In this setup, the estimator of Eq. (27) can be re-expressed as:
$\hat{I}^{\rm(QAIS,obs)}_{N}=\frac{1}{N}\sum_{i\in\Omega^{-}}w^{(i)}\sum_{j=1}^{N_{i}}f({\bf x}^{(i)}_{j}).$ (29)
The missing contribution from $\Omega\setminus\Omega^{-}$ introduces a systematic bias to the integral estimation, leading to an underestimation in its absolute value. This bias can be explicitly evaluated as
$\text{Bias}_{\Omega\setminus\Omega^{-}}=\mathbb{E}\left[\hat{I}^{\rm(QAIS,obs)}_{N}\right]-I=-\sum_{i\in\Omega\setminus\Omega^{-}}I_{i}~,$ (30)
@@ 3.3 Debiasing Strategy using a Tiling Algorithm
$\hat{I}^{\rm(QAIS)}_{N}=\frac{1}{N}\sum_{i\in\Omega_{I}}w^{(i)}\sum_{j=1}^{N_{i}}f({\bf x}^{(i)}_{j})+\frac{1}{N}\sum_{k\in\Omega_{N-I}}\tilde{w}^{(k)}\sum_{j=1}^{\tilde{N}_{k}}f({\bf x}^{(k)}_{j})~,$ (31)
where the adjusted weights $\tilde{w}^{(k)}$ take into account the shapes of the new enlarged hyper-rectangular domains and $\tilde{N}_{k}$ refers to the number of random samples drawn within each enlarged domain $\tilde{\Omega}_{k}$. Because the debiasing step essentially groups the cells, re-labels them and repositions some of the existing $N$ samples, the estimator’s formula of Eq. (31) is equivalent to Eq. (27), with the variance given by Eq. (28), provided that the adjusted weights are used where it is appropriate. Through this construction, we explicitly include all previously neglected states, eliminating missing regions and ensuring that the estimator is unbiased.
In practice, for constructing the non-Important Region, we employ a Tiling algorithm. This algorithm generates a small number of contiguous, non-overlapping hyper-rectangles that collectively cover every previously unobserved grid cell, a process that corresponds to a full tiling of the $d$-dimensional integration domain. For the construction of such an algorithm, it is important to have a computational cost that scales polynomially with the number of observed states $M$, the dimension of the integration space $d$, and the number of qubits used for discretization $n$. It is an explicit and crucial constraint for the Tiling algorithm not to scale with the number of unobserved grid cells, since, in this framework, this is an exponential sized quantity.
@@
[…]
@@ 3.4 Structure and Optimization of the Parametrized Quantum Circuit
To convert the continuous function $f$, into a discrete PDF or a Probability Mass Function (PMF) suitable for computing the discretized KL-Divergence, we proceed as follows. We first compute the function at $N_{i}$ representative points $\mathbf{x}_{i}$} within each observed grid cell $\Omega^{(i)}$, accounting in total to $N=\sum_{i=1}^{M}N_{i}$ function evaluations. Because the grid is uniform, every cell has the same volume, so common factors cancel in the normalization. Therefore, we construct the PMF by taking the simplified expressions, using $P(\Omega^{(i)})=f(\Omega^{(i)})/Z$ which is the probability of each grid cell, where $f(\Omega^{(i)})=1/N_{i}\sum_{i=1}^{N_{i}}f(\mathbf{x}_{i})$ is the within-cell function’s average and $Z=\sum_{i=1}^{M}f(\Omega^{(i)})$ the normalization constant. For the non-uniform grid case, the cell’s estimate is computed as $f(\Omega^{(i)})=|\Omega^{(i)}|/N_{i}\sum_{i=1}^{N_{i}}f(\mathbf{x}_{i})$, with the resulting change in $Z$. Ideally, for a better precision in the target PMF, the cell-wise sample count, $N_{i}$ would be updated after each run, by assigning more sample to high probability cells. This, combined with the monitoring of the within-cell variance in high impact region, can provide a criterion that indicates whether the grid resolution must be increased by adding more qubits. However, in this work, for robustness and simplicity, we use a fixed $N_{i}$ per cell and cache all evaluations, while pre-defining the grid’s size.
@@
Figure 2: Quantum Circuit architecture used for training. The figure features the general structure with the layer definition (top) as well as an explicit 555 qubit example (bottom).
@@ 3.5 Absence of phantom peaks
where $\mathbf{r}_{0}=(0.23,0.23)$ and $\mathbf{r}_{1}=(0.74,0.74)$. Because the integral can be computed in closed form, it provides a clear benchmark for validating numerical methods. A normalization constant is chosen via analytical integration, so that the integral evaluates to 111.
@@
To demonstrate that QAIS is unbiased, we conducted a series of nested MC experiments. Because the integral is analytically normalized to 1, this exact value serves as the ground truth for validating our numerical estimates. For each experiment, we performed 100010001000 independent integration runs, each with a fixed proposal PDF from the same trained PQC. We then repeated these experiments with different shot counts: 100010001000, 500050005000 and 100001000010000 samples. The results are shown in Fig. 4. We see that the average integral value oscillates around the true value, improving accuracy progressively, and does not exhibit a systematic preference toward over or underestimation when the number of samples is varied. Additionally, the average QAIS standard deviation is presented. It shall be noted that estimates that deviate significantly from the mean integral value, always exhibit inflated standard deviation. Thus, QAIS produced accurate and unbiased estimates without the need to sample all $2^{n}$ possible states.
@@
where $\mathbf{r}=(0.5,0.5)$. Figure 5 illustrates the results for both methods. The exponential factor puts almost the entire weight of the integrand on a thin ring of radius 0.350.350.35 centered at $\mathbf{r}$, while in the rest of the integration domain, it is exponentially suppressed and effectively negligible.
@@ 4.1 One-loop pentagon Feynman integral in the Loop-Tree Duality
The QAIS results are presented in Fig. 7 and Fig. 8. The training of the PQC was carried out with COBYLA, the Pennylane’s lightning.qubit noiseless state‑vector simulator, and different Anzätze with different numbers of qubits per dimension. It is worth noting that, in the smallest configuration we tested, i.e. the 16 qubits distributed as (8,4,4)(8,4,4)( 8 , 4 , 4 ) across the three dimensions and $2\times 10^{4}$ iterations, the Ansatz sequence $U^{(Z)}\to U_{3}$ with 144144144 parameters reaches $D_{\rm KL}<0.5$ very fast and then saturates at $D_{\rm KL}\approx 0.27$. The results presented in Fig. 7 are obtained using the Ansatz $U^{(Z)}\to U_{3}\to U^{(X)}\to U_{3}$, which performed best given the constraints we imposed on the number of iterations. In this case, the number of parameters ranges from 336336336 to 456456456 for 161616 to 191919 qubits, respectively, and the number of iterations ranges from $2\times 10^{4}$ to $3\times 10^{4}$. The $D_{\rm KL}$ ranged from 0.090.090.09 to 0.140.140.14.
For the HEA, with the current number of iterations for the (8,4,4)(8,4,4)( 8 , 4 , 4 ) qubit case, we tested between 222 to 888 layers, corresponding to 969696 to 384384384 parameters, with different random initializations. The results obtained were at least one order of magnitude larger in the KL divergence ($D_{\rm KL}>1$), indicating a significant discrepancy. It should be noted that the optimization that starts from a uniform superposition has initial $D_{\rm KL}\approx 6.2$.
@@
Figure 8: Histograms of the values of the 100 independent integration runs for the one-loop pentagon Feynman diagram with parameters in Eq. (40). The dashed vertical lines, correspond to $1\sigma$, $2\sigma$, and $3\sigma$ standard deviations. The reference value comes by the comparison with the classical numerical integration using VEGAS Buchta:2015wna , that is $I=-1.24027(16)\cdot 10^{-13}$.